# Ultra-sensitive sizing of individual nanoparticles with an optofluidic microcavity

### v1

@@ Abstract
Nanomaterials are often heterogeneous, and methods that give direct access to single particle properties can open up the view into the details of the distribution, reveal features that are otherwise washed out in ensembles, and enable the studies of dynamical behaviour. Therefore, a range of label-free single-particle sensing techniques has been developed. For example, interferometric scattering microscopy (iSCAT) [1, 2] has enabled the detection of single nanoparticles and biomolecules [3] with sizes below 10 kDa using machine learning algorithms [4], and the tracking [5] or sizing [6] of single nanoparticles down to 10 nm in diameter.
Another approach is the use of nanoplasmonic hot spots, e.g. at the tip of gold nanorods [7, 8], or whispering gallery mode (WGM) resonators with high quality factors such as microspheres [9, 10], which enable real-time detection of single particles as well as quantitative particle sizing [11]. Very high sensitivity is achievable by combining WGM cavities with nanoplasmonic hot spots [12]. However, most of these sensors are based on the interaction of the analyte with the near-field and thus the sensor surface, which may constrain and modify diffusion. This can preclude a precise quantification of, for example, the hydrodynamic radius. Open access microcavities avoid this limitation [13, 14, 15, 16], but the larger mode volume compared to nanoplasmonic sensors has so far restricted detection and sizing to particles $>~{}100$ nm. Only recently, open cavities could achieve sensitivities for single molecules down to $\sim 1$ kDa using a highly non-linear photothermal regime [17]. Although this remarkable sensitivity has allowed for single protein detection, extracting information relating to the size or shape with this approach has remained limited to the statistics obtained from many detection events. In addition, the influence of the photothermal response that convolves with the signal makes quantitative and unbiased evaluation challenging.
@@
While such an analysis is well established for ensemble techniques such as fluorescence correlation spectroscopy (FCS) and dynamic light scattering (DLS) [18], the scenario studied here raises two central aspects that require careful attention: First, the probing volume in FCS is well approximated by a 3D Gaussian that can be treated with an analytical model for the ACF, while a cavity standing wave has a more complex structure that requires a reformulation of this model. Second, the extraction of diffusion time constants and hydrodynamic radii from the ACF relies on the assumption of a stationary system, which can be satisfied by an ensemble with a constant average concentration. In contrast, single particle events are intrinsically non-stationary. We address both aspects and propose an analytical model for the ACF of particle diffusion through a standing wave probe volume, and introduce an analysis that minimizes undesired influence of the transient single-particle behavior. With this methodology, we perform quantitative particle sizing for a range of nominal particle diameters and obtain size distributions that are in good agreement with reference measurements performed with DLS and transmission electron microscopy (TEM). This opens the way to quantitatively assess nanomaterials such as unlabeled biomolecules in their native environment, and could allow further dynamic processes such as rotation and conformational changes to be investigated in the future.
@@
I Results
I.1 The locked optofluidic microcavity
Fabry-Perot microcavities with high finesse strongly enhance the interaction between light and matter in the cavity mode volume, and therefore are powerful tools for the label-free detection of single nanoparticles in aqueous dispersion. Similarly to previous work [15], we use a fiber-based optofluidic microcavity comprising two single-mode optical fibers inserted into a fiber-optic ferrule, with a microfluidic channel drilled into the ferrule perpendicular to the fibers, such that the cavity between the two fiber tips is immersed in water, see Fig. 1 (a, b). The optical fibers are processed by laser machining to produce concave profiles with radii of curvature between 20 µm and 80 µm, and then coated with a highly reflective Bragg mirror having transmission $\mathcal{T}=$ 20 ppm in air [19].
We implement an actively stabilized cavity locked on the side of a resonance using a side-of-fringe-lock with an FPGA-based feedback circuit with a unity-gain bandwidth of 80 Hz, which is well below the frequency range covered by nanoparticle diffusion. In comparison to previously reported length-modulated cavities [15, 13], the locked cavity offers the possibility of measurements at much higher bandwidth, limited essentially only by the cavity linewidth.
The cavity lengths used in our experiments were 3.7 – 4.5 µm and the cavity finesse in water was ${4.5\times 10^{4}}-{5.8\times 10^{4}}$. Under active stabilization, the microcavity shows a root-mean-square optical length instability of $\sim$ 300 fm, which corresponds to about 5% of the linewidth, over several hours. The cavity-locked transmission trace is illustrated in Fig. 1. We note that even the passively stable cavity shows high stability on short timescales (up to seconds), providing a low noise background for the measurements.
I.2 Detection of single nanoparticles
@@
We use the microcavity to detect single citrate-stabilized spherical gold nanoparticles (GNP) in aqueous dispersion with specified diameters ranging from 3 nm to 20 nm. The stock sample was diluted to a concentration of 50 pM, so that there is, on average, less than one particle in the cavity mode at any time. The resulting suspension is injected into the microcavity, and the cavity transmission $y(t)$ is monitored with an avalanche photodiode on a 12-bit digital storage oscilloscope. Examples of single nanoparticle events are shown in Fig. 2 (a) and the top panels of (c) and (d).
@@
Figure 2: (a) Example of the transmission of a locked cavity disturbed by the diffusion of a 20 nm nanosphere through the optical mode. (b) The autocorrelation of the nanoparticle signal and its fit to the analytical ACF presented in this work. (c)–(d) Time traces (top panel), correlation amplitudes $G(\delta\tau)$ (middle panel), and autocorrelograms (bottom panel) of 20 nm and 10 nm nanosphere events respectively. (e) Each nanoparticle can be mapped onto the two-dimensional space of nanoparticle diameter and autocorrelation amplitude contrast $\Delta G(\delta\tau)$.
@@
We calculate the ACF $G(\tau)$ of the measured signal and observe characteristic decays of the correlation on two separated time scales, see Fig. 2 (b). This differs from the autocorrelation typically observed by other methods when only translational diffusion is present [20], and we now analyze and model this behavior in detail to enable a fit of the data and extract physical quantities of interest, such as the hydrodynamic radius. As a first step, we use the autocorrelation to enable the detection of nanoparticles, even when the signal-to-noise ratio of the measured cavity transmission time trace does not allow a direct identification of nanoparticle events. For this purpose, we compute the autocorrelation of short segments of the data with duration $T_{\mathrm{seg}}$ = 10 ms, and evaluate the correlation amplitude $G(\delta\tau)$, where $\delta\tau$ is the shortest temporal delay in the autocorrelation, for each segment. The middle panels of Fig. 2 (c) and (d) show examples of $G(\delta\tau)$ for 20 nm and 10 nm particles, respectively, while the lower panel shows the merging of ACFs of the segments into a time-resolved autocorrelogram. Whenever a particle is present, $G(\delta\tau)$ shows a value significantly larger than without a particle, since the empty-cavity background signal is largely uncorrelated. We thus use the autocorrelation amplitude contrast $\Delta G(\delta\tau)=G_{\mathrm{NP}}(\delta\tau)-G_{\mathrm{BG}}(\delta\tau)$ between the peak correlation amplitude $G_{\mathrm{NP}}(\delta\tau)$ of such a segment and the background correlation amplitude, $G_{\mathrm{BG}}(\delta\tau)$, as a measure of particle presence. The statistical distribution of $\Delta G(\delta\tau)$ for GNP events with different particle sizes is shown in Fig. 2 (e), where a roughly exponential correlation of $\Delta G(\delta\tau)$ with particle size can be observed. This shows that $\Delta G(\tau)$ alone already allows to differentiate particles of different size, but only qualitatively and with comparably large uncertainty. We note that the background value $G_{\mathrm{BG}}(\delta\tau)$ for times where no particle is present changes slightly between measurements due to changing noise levels (see Supplementary Information). However, the background remains a factor of $\sim 2$ smaller even for 3 nm GNPs. The use of the signal correlation contrast is thus a powerful way to detect very small particles whose signal would otherwise remain masked by noise. While it enables highly sensitive particle detection, provides qualitative size information, and does not require any modeling of the ACF, it does not provide a quantitative characterization of the particles. We therefore proceed to develop a proper model to enable particle sizing as well.
I.3 The autocorrelation function of diffusion in a standing wave field
@@
Autocorrelation-based analysis of diffusional motion is a widely used technique to determine the size of particles in suspensions by analyzing their diffusional motion, for instance in FCS [20, 21, 22]. For the accurate determination of time constants of diffusion, it is imperative to first determine the theoretical ACF for the optical mode geometry involved.
@@
where $x=1+2D_{T}k^{2}s\tau$, and $a_{m},b_{m}$ are analytically determined constants. In Fig. 3 (c), we compare the ACFs arising from different numbers of Gaussian antinodes used. We note that when $q=1$, the ACF corresponds to a single three-dimensional Gaussian focus and is identical to that used in FCS (Eqn. 1). When $q>1$, a second decay appears, which arises from diffusion through to multiple antinodes, i.e. due to the periodicity of the field seen by the nanoparticle. Additionally, the ACF converges as $q\gg 1$, justifying the simplification we used to obtain Eqn. 3. Furthermore, the transverse dimension of the cavity mode leads to a third decay time constant, shown as a grey dotted line in Fig. 3. However, it overlaps with the multiple-antinode decay for our cavity geometry, and cannot be resolved.
@@
I.4 Autocorrelation of single particle events
@@
We now turn to the aspect of treating a transient signal with possible contributions from the lock, which requires additional considerations and signal processing. In the first step, the offset of the lock set point and the influence of small external disturbances or of the locking mechanism in long particle events are removed by baseline-correcting the particle signal (see Supplementary Information for further details).
@@
Second, a single nanoparticle event occupies only a part of a typical measurement, and we wish to select only this segment for calculation of the ACF. For larger particles, this region of interest can be identified by monitoring the amplitude of fluctuations above the background noise level; however, for smaller particles ($R<$ 10 nm), this is not reliable due to the small signal-to-noise ratio they produce. Therefore, we calculate $G(\delta\tau)$ as introduced above and use this value as a measure of particle presence in the trace segment, similar to [23]. Segments with a low correlation amplitude are excluded from further analysis. It can be seen from Fig. 2 (c) and (d) that segments exhibiting higher correlation than the background can thus be effectively identified and extracted for further processing. This method for automatic nanoparticle recognition reduces bias towards larger particles or agglomerates compared to amplitude-based recognition, and allows the extraction of particle signals with fluctuation amplitudes comparable to the noise background.
@@
Third, having identified a nanoparticle event in the measured signal, the transient nature of a single-particle event must be taken into account. The analysis of a fluctuation signal as described in Sec. I.3 is valid for a stationary system, in which the fluctuations in particle concentration are small compared to the average concentration in the observation volume. In the regime of a single-particle measurement, however, the average concentration of particles in the observation volume lies in ${0<\langle C(t)\rangle<1}$, depending on the individual particle trajectory. As a result, one cannot arbitrarily subtract the mean signal level to obtain the fluctuation signal. Instead, we note from Fig. 3 (c) that the ACF derived for a standing wave shows an initial decay corresponding to diffusion through a single antinode of the optical field and a second decay corresponding to motion between antinodes, separated by a point of minimum slope. We find that the height of the point of minimum slope depends only on the cavity geometry and not on the sample. We can thus use this point as a reference level to which the amplitude of the first diffusion decay is normalized. This is done by subtracting an offset $y_{0}$ from the measured signal time trace, which effectively corresponds to $\langle F(t)\rangle$ in Sec. I.3. In Fig. 4 (a)–(c), we show measured ACFs of single particle events from particles with nominal diameters of 20 nm, 15 nm, and 10 nm, respectively.
@@
We fit the thus obtained experimental ACFs with the analytical model introduced above over the extent of the first diffusion decay (Fig. 2 (b)). The second decay is not useful due to the limited statistics it represents. We use $D_{T}$ as the only fit parameter, which is directly related to the hydrodynamic radius $R$ via the Stokes-Einstein relation. Hence, we can extract the size of the single particle from each event. Exemplary fits are shown for the average ACFs, and a fit to a single particle signal is shown in Fig. 2 (b).
@@
To verify the accuracy of our method, we compare our results to measurements with a commercial DLS system and TEM images, which are in excellent agreement as shown in Fig. 4 (d), (e). The deviations of the measured sizes from the nominal values points towards systematic uncertainties in the nominal values given by the manufacturer. It can be observed that for the smallest two particle sizes, the cavity measurement reproduces the right size trend and is very close to the nominal diameter, while DLS shows an opposing trend, pointing towards the limitation of the latter method. Also, we note an intrinsic ambiguity in DLS results, since the scattered intensity due to a particle with radius $R$ in an ensemble measurement is proportional to $R^{6}$. As a result, the directly measured intensity-normalized distributions are highly vulnerable to subpopulations of larger particles or agglomerates. Instead, we have considered number-normalized distributions which, while offering a more meaningful comparison in this work, are calculated using assumptions about some nanoparticle properties. They are thus dependent on a priori knowledge of the nanoparticles and are susceptible to errors in the provided data. Finally, in contrast to DLS measurements, only microscopic sample volumes of a few hundred microliters with ultra-low concentration are required for the microcavity experiments.
II Conclusions

### v2

# Accurate single-nanoparticle sizing down to 3 nm with an optofluidic microcavity

@@ Abstract
Keywords: single-particle detection, Brownian motion, autocorrelation, sizing, optical microcavity, gold nanoparticles.
[…]
Karlsruher Institut für Technologie, Institut für QuantenMaterialien und Technologien , Hermann-von-Helmholtz-Platz 1, 76344 Eggenstein-Leopoldshafen, Germany
1 Introduction
Nanomaterials are often heterogeneous, and methods that give direct access to single particle properties can open up the view into the details of the distribution, reveal features that are otherwise washed out in ensembles, and enable the studies of dynamical behaviour. Therefore, a range of label-free single-particle sensing techniques has been developed. For example, interferometric scattering microscopy (iSCAT) 1, 2 has enabled the detection of single nanoparticles and biomolecules 3 with sizes below 10 kDa using machine learning algorithms 4, and the tracking 5 or sizing 6 of single nanoparticles down to 10 nm in diameter.
Another approach is the use of nanoplasmonic hot spots, e.g. at the tip of gold nanorods 7, 8, or whispering gallery mode (WGM) resonators with high quality factors such as microspheres 9, 10, which enable real-time detection of single particles as well as quantitative particle sizing 11. Very high sensitivity is achievable by combining WGM cavities with nanoplasmonic hot spots 12. However, most of these sensors are based on the interaction of the analyte with the near-field and thus the sensor surface, which may constrain and modify diffusion. This can preclude a precise quantification of, for example, the hydrodynamic radius. Open access microcavities avoid this limitation 13, 14, 15, 16, but the larger mode volume compared to nanoplasmonic sensors has so far restricted detection and sizing to particles >100>~100 nm. Only recently, open cavities could achieve sensitivities for single molecules down to $\sim 1$ kDa using a highly non-linear photothermal regime 17. Although this remarkable sensitivity has allowed for single protein detection, extracting information relating to the size or shape with this approach has remained limited to the statistics obtained from many detection events. In addition, the influence of the photothermal response that convolves with the signal makes quantitative and unbiased evaluation challenging.
@@ 1 Introduction
While such an analysis is well established for ensemble techniques such as fluorescence correlation spectroscopy (FCS) and dynamic light scattering (DLS) 18, the scenario studied here raises two central aspects that require careful attention: First, the probing volume in FCS is well approximated by a 3D Gaussian that can be treated with an analytical model for the ACF, while a cavity standing wave has a more complex structure that requires a reformulation of this model. Second, the extraction of diffusion time constants and hydrodynamic radii from the ACF relies on the assumption of a stationary system, which can be satisfied by an ensemble with a constant average concentration. In contrast, single particle events are intrinsically non-stationary. We address both aspects and propose an analytical model for the ACF of particle diffusion through a standing wave probe volume, and introduce an analysis that minimizes undesired influence of the transient single-particle behavior. With this methodology, we perform quantitative particle sizing for a range of nominal particle diameters and obtain size distributions that are in good agreement with reference measurements performed with DLS and transmission electron microscopy (TEM). This opens the way to quantitatively assess nanomaterials such as unlabeled biomolecules in their native environment, and could allow further dynamic processes such as rotation and conformational changes to be investigated in the future.
@@
2 Results
2.1 The locked optofluidic microcavity
Fabry-Perot microcavities with high finesse strongly enhance the interaction between light and matter in the cavity mode volume, and therefore are powerful tools for the label-free detection of single nanoparticles in aqueous dispersion. Similarly to previous work 15, we use a fiber-based optofluidic microcavity comprising two single-mode optical fibers inserted into a fiber-optic ferrule, with a microfluidic channel drilled into the ferrule perpendicular to the fibers, such that the cavity between the two fiber tips is immersed in water, see Fig. 1 (a), (b). The optical fibers are processed by laser machining to produce concave profiles with radii of curvature between 20 µm and 80 µm, and then coated with a highly reflective Bragg mirror having transmission $\mathcal{T}=$ 20 ppm in air 19.
We implement an actively stabilized cavity locked on the side of a resonance using a side-of-fringe-lock with an FPGA-based feedback circuit with a unity-gain bandwidth of 80 Hz, which is well below the frequency range covered by nanoparticle diffusion (see Supporting Information, Figs. S3 and S4). In comparison to previously reported length-modulated cavities 15, 13, the locked cavity offers the possibility of measurements at much higher bandwidth, limited essentially only by the cavity linewidth.
The cavity lengths used in our experiments were 3.7 – 4.5 µm and the cavity finesse in water was ${4.5\times 10^{4}}-{5.8\times 10^{4}}$. Under active stabilization, the microcavity shows a root-mean-square optical length instability of $\sim$ 300 fm, which corresponds to about 5% of the linewidth, over several hours. The intracavity power circulating in the locked cavity is in the order of 50 mW. A cavity-locked transmission trace is illustrated in Fig. 1. We note that even the passively stable cavity shows high stability on short timescales (up to seconds), providing a low noise background for the measurements.
2.2 Detection of single nanoparticles
@@
We use the microcavity to detect single citrate-stabilized spherical gold nanoparticles (GNP) in aqueous dispersion with specified diameters ranging from 3 nm to 20 nm. The stock sample was diluted to a concentration of 50 pM, so that there is, on average, less than one particle in the cavity mode at any time. The resulting suspension is injected into the microcavity, and the cavity transmission $y(t)$ is monitored with an avalanche photodiode on a 12-bit digital storage oscilloscope. Examples of single nanoparticle events are shown in Fig. 2 (a) and the top panels of (c) and (d), as well as Fig. S5 in the Supporting Information.
@@
Figure 2: (a) Example of the transmission of a locked cavity disturbed by the diffusion of a 20 nm nanosphere through the optical mode. (b) The autocorrelation of the nanoparticle signal and its fit to the analytical ACF developed in this work, along with the relatively flat background autocorrelation. (c)–(d) Time traces (top panel), correlation amplitudes $G(\delta\tau)$ (middle panel), and autocorrelograms (bottom panel) of 20 nm and 10 nm nanosphere events respectively. (e) Each nanoparticle can be mapped onto the two-dimensional space of nanoparticle diameter and autocorrelation amplitude contrast $\Delta G(\delta\tau)$.
@@
We calculate the ACF $G(\tau)$ of the measured signal and observe characteristic decays of the correlation on two separated time scales, see Fig. 2 (b). This differs from the autocorrelation typically observed by other methods when only translational diffusion is present 20, and we now analyze and model this behavior in detail to enable a fit of the data and extract physical quantities of interest, such as the hydrodynamic radius. As a first step, we use the autocorrelation to enable the detection of nanoparticles, even when the signal-to-noise ratio of the measured cavity transmission time trace does not allow a direct identification of nanoparticle events. For this purpose, we compute the autocorrelation of short segments of the data with duration $T_{\mathrm{seg}}$ = 10 ms, and evaluate the correlation amplitude $G(\delta\tau)$, where $\delta\tau$ is the shortest temporal delay in the autocorrelation, for each segment. The middle panels of Fig. 2 (c) and (d) show examples of $G(\delta\tau)$ for 20 nm and 10 nm particles, respectively, while the lower panel shows the merging of ACFs of the segments into a time-resolved autocorrelogram. Whenever a particle is present, $G(\delta\tau)$ shows a value significantly larger than without a particle, since the empty-cavity background signal is largely uncorrelated. We thus use the autocorrelation amplitude contrast $\Delta G(\delta\tau)=G_{\mathrm{NP}}(\delta\tau)-G_{\mathrm{BG}}(\delta\tau)$ between the peak correlation amplitude $G_{\mathrm{NP}}(\delta\tau)$ of such a segment and the background correlation amplitude, $G_{\mathrm{BG}}(\delta\tau)$, as a measure of particle presence. The statistical distribution of $\Delta G(\delta\tau)$ for GNP events with different particle sizes is shown in Fig. 2 (e), where a roughly exponential correlation of $\Delta G(\delta\tau)$ with particle size can be observed. This shows that $\Delta G(\delta\tau)$ alone already allows to differentiate particles of different size, but only qualitatively and with comparably large uncertainty. We note that the background value $G_{\mathrm{BG}}(\delta\tau)$ for times where no particle is present changes slightly between measurements due to changing noise levels (see Supporting Information, Fig. S5 (f)). However, the background remains a factor of $\sim 2$ smaller even for 3 nm GNPs. The use of the signal correlation contrast is thus a powerful way to detect very small particles whose signal would otherwise remain masked by noise. While it enables highly sensitive particle detection, provides qualitative size information, and does not require any modeling of the ACF, it does not provide a quantitative characterization of the particles. In addition, the magnitude of $\Delta G(\delta\tau)$ is dependent on parameters of our measurement system, such as cavity geometry, measurement bandwidth and the background, and hence does not allow the determination of nanoparticle properties without a priori knowledge or calibration measurements. We therefore proceed to develop a proper model to enable accurate particle sizing.
2.3 The autocorrelation function of diffusion in a standing wave field
@@
Autocorrelation-based analysis of diffusional motion is a widely used technique to determine the size of particles in suspensions by analyzing their diffusional motion, for instance in FCS 20, 21, 22. For the accurate determination of time constants of diffusion, it is imperative to first determine the theoretical ACF for the optical mode geometry involved.
@@
where $x=1+2D_{T}k^{2}s\tau$, and $a_{m},b_{m}$ are analytically determined constants. In Fig. 3 (c), we compare the ACFs arising from different numbers of Gaussian antinodes used. We note that when $q=1$, the ACF corresponds to a single three-dimensional Gaussian focus and is identical to that used in FCS (Eqn. 1). When $q>1$, a second decay appears, which arises from diffusion through to multiple antinodes, i.e. due to the periodicity of the field seen by the nanoparticle. Additionally, the ACF converges as $q\gg 1$, justifying the simplification we used to obtain Eqn. 3. Furthermore, the transverse dimension of the cavity mode leads to a third decay time constant, shown as a grey dotted line in Fig. 3. However, it overlaps with the multiple-antinode decay for our cavity geometry, and cannot be resolved. We verify that the approximations introduced do not cause systematic errors in subsequent ACF evaluation by Monte Carlo simulations (see Supporting Information, Fig. S1).
@@
2.4 Autocorrelation of single particle events
@@
We now turn to the aspect of treating a transient signal with possible contributions from the lock, which requires additional considerations and signal processing.
In the first step, the offset of the lock set point and the influence of small external disturbances or of the locking mechanism in long particle events are removed by baseline-correcting the particle signal (see Supporting Information).
@@
Second, a single nanoparticle event occupies only a part of a typical measurement, and we wish to select only this segment for calculation of the ACF. For larger particles, this region of interest can be identified by monitoring the amplitude of fluctuations above the background noise level; however, for smaller particles ($R<$ 10 nm), this is not reliable due to the small signal-to-noise ratio they produce. Therefore, we calculate $G(\delta\tau)$ as introduced above and use this value as a measure of particle presence in the trace segment, similar to the approach of Asgari et al 23. Segments with a low correlation amplitude are excluded from further analysis. It can be seen from Fig. 2 (c) and (d) that segments exhibiting higher correlation than the background can thus be effectively identified and extracted for further processing. This method for automatic nanoparticle recognition reduces bias towards larger particles or agglomerates compared to amplitude-based recognition, and allows the extraction of particle signals with fluctuation amplitudes comparable to the noise background.
@@
Third, having identified a nanoparticle event in the measured signal, the transient nature of a single-particle event must be taken into account. The analysis of a fluctuation signal as described earlier is valid for a stationary system, in which the fluctuations in particle concentration are small compared to the average concentration in the observation volume. In the regime of a single-particle measurement, however, the average concentration of particles in the observation volume lies in $0<\langle C(t)\rangle<1$}, depending on the individual particle trajectory. As a result, one cannot arbitrarily subtract the mean signal level to obtain the fluctuation signal. Instead, we note from Fig. 3 (c) that the ACF derived for a standing wave shows an initial decay corresponding to diffusion through a single antinode of the optical field and a second decay corresponding to motion between antinodes, separated by a point of minimum slope. We empirically find that the height of the point of minimum slope depends only on the cavity geometry and not on the sample. We can thus use this point as a reference level to which the amplitude of the first diffusion decay is normalized. This is done by subtracting an offset $y_{0}$ from the measured signal time trace, which effectively corresponds to $\langle F(t)\rangle$.
@@
In Fig. 4 (a)–(c), we show measured ACFs of single particle events from particles with nominal diameters of 20 nm, 15 nm, and 10 nm, respectively. We fit the thus obtained experimental ACFs with the analytical model introduced above over the extent of the first diffusion decay (Fig. 2 (b)). The second decay is less reliable due to the smaller statistics it represents. We use $D_{T}$ as the only fit parameter, which is directly related to the hydrodynamic radius $R$ via the Stokes-Einstein relation. Hence, we can extract the size of the single particle from each event. Exemplary fits are shown for the average ACFs, and a fit to a single particle signal is shown in Fig. 2 (b). We emphasize the importance of our careful treatment by noting that using the ACF for a single Gaussian as in FCS would lead to systematic errors and increased uncertainty (see Supporting Information, Fig. S2 and Table S1).
@@
To verify the accuracy of our method, we compare our results to measurements with a commercial DLS system and TEM images, which are in excellent agreement as shown in Fig. 4 (d), (e). The deviations of the measured sizes from the nominal values point towards systematic uncertainties in the nominal values given by the manufacturer. It can also be observed that for the smallest two particle sizes, the cavity measurement reproduces the right size trend and is very close to the nominal diameter, while DLS shows an opposing trend, pointing towards the limitation of the latter method. Also, we note an intrinsic ambiguity in DLS results, since the scattered intensity due to a particle with radius $R$ in an ensemble measurement is proportional to $R^{6}$. As a result, the directly measured intensity-normalized distributions are highly vulnerable to subpopulations of larger particles or agglomerates. Instead, we have considered number-normalized distributions which, while offering a more meaningful comparison in this work, are calculated using assumptions about some nanoparticle properties. They are thus dependent on a priori knowledge of the nanoparticles and are susceptible to errors in the provided data. Finally, in contrast to DLS measurements, only microscopic sample volumes of a few hundred microliters with ultra-low concentration are required for the microcavity experiments.
With the TEM measurements of $N=34$ and 20 individual GNPs of size 10 nm and 20 nm respectively, we compare our results to another single-particle technique, again with very good agreement, as shown in Fig. 4 (i). Notably, the width of the size distribution obtained from cavity measurements matches the one from TEM measurements and is more narrow than the one from DLS. Also, the mean values of the particle size agree very closely. This indicates a high precision and accuracy of our method. Other precision single-particle sizing measurements in the literature, such as nanoparticle tracking analysis (NTA) and interferometric NTA (iNTA) experiments, result in similar distributions. 6 We note, however, that our measurements enable sizing of more than three times smaller particles.
3 Conclusions